Derive the expression for electric potential due a uniformly charged spherical
shell at a point outside the shell.
1
Expert's answer
2020-04-21T07:57:02-0400

Let, R be the radius of the shell and Q be the charge uniformly distributed on the surface.

i) For a point outside the shell :

By Gauss's law, "E.4\u03c0r ^\n2\n =\n\u200b\t\n \nQ _{\nen}\/\\epsilon_0\n\n\u200b"

here r be the distance from centre of shell (r>R) and charge enclosed by surface "Qeq\n\n\u200b=Q."

So, the electic field outside the shell is  "E\n =\n\u200b\t\n \nQ _{\nen}\/(4\u03c0r^2\\epsilon_0)" , the electric potential outside the shell is "V(r)=-\int_{\infin}^r \frac{Q_{en}}{4\pi r_1^2 \epsilon_0}d r_1=\frac{Q_{en}}{4\pi r\epsilon_0}."